### v1

# Nonreciprocal superradiant quantum phase transition induced by magnon Kerr effect

Abstract
Recently, proposals for realizing a nonreciprocal superradiant quantum phase transition (SQPT) have been put forward, based on either nonreciprocal interactions between two spin ensembles or the Sagnac-Fizeau shift in a spinning cavity. However, experimental implementation of such a nonreciprocal SQPT remains challenging. This motivates the search for new mechanisms capable of producing a nonreciprocal SQPT. Here, we propose an alternative approach to realize a nonreciprocal SQPT, induced by the magnon Kerr effect (MKE), in a cavity magnonic system, where magnons in a yttrium iron garnet (YIG) sphere are coupled to cavity photons. The Kerr coefficient $K$ is positive (negative), i.e., $K>0$ ($K<0$), when the bias magnetic field is aligned along the crystallographic axis [100] ([110]) of the YIG sphere. We show that the steady-state phase diagrams differ significantly for both cases with $K>0$ and $K<0$, which is the origin of the nonreciprocal SQPT. By further studying the steady-state magnon occupation and its fluctuations versus the parametric drive strength, we demonstrate that the SQPT becomes nonreciprocal, characterized by distinct critical thresholds for $K>0$ and $K<0$. Moreover, we introduce a bidirectional contrast ratio to quantify this nonreciprocal behavior. Our work provides a new mechanism for realizing the nonreciprocal SQPT, with potential applications in designing nonreciprocal quantum devices.
I Introduction
Superradiant quantum phase transition (SQPT) in the Dicke model was first predicted by Hepp et al. in 1973 Hepp73 ; Wang73 . The Dicke model describes the collective coupling of $N$ two-level atoms to a quantized light field Dicke54 . In the thermodynamic limit with the number of atoms $N\rightarrow+\infty$, the system undergoes a SQPT from the normal phase to the superradiant phase at zero temperature as the coupling strength exceeds a critical value Emary03PRL ; Emary03PRE . Below the critical coupling, the system is in the normal phase, where the atoms remain in their ground state and the cavity field is in the vacuum state. Above the critical coupling, the system enters the superradiant phase, characterized by collective excitations of both the atomic ensemble and the cavity field. With advances in experimental capabilities, the SQPT in the Dicke model and its variants (e.g., Tavis-Cummings model Castanos09 ; Zou13 , Rabi model Ashhab13 ; Hwang15 , Jaynes-Cummings model Hwang16 ; Liu24 , and coupled multi-oscillator model Zhang21-Chen ; Qin22 ; Wang24Nori ) has been observed in a variety of quantum platforms. These include Bose-Einstein condensates within optical cavities Baumann10 ; Baumann11 ; Brennecke13 ; Baden14 , degenerate Fermi gases inside an optical cavity Zhang21Chen ; Wu23 , circuit QED systems Feng15 ; Zheng23 ; Ning24 , trapped ions Cai21 ; Zhao25 , nuclear-magnetic-resonance systems Chen21 ; Wu24 , and so on.
Recently, Fruchart et al. established a general theory of nonreciprocal phase transitions, showing that time-dependent phases arise from asymmetric interactions among multiple species or fields Fruchart21 . The nonreciprocal phenomenon refers to that the system responds differently when an external field is applied in opposite directions. Due to potential applications in quantum technologies, various nonreciprocal quantum effects have been studied, such as thermal noise routers Barzanjeh18 , nonreciprocal photon blockade Huang18 , one-way quantum amplifiers Metelmann15 ; Malz18 ; Shen18 , and nonreciprocal magnon blockade Wang22 . Following the framework of Ref. Fruchart21 , the nonreciprocal SQPT was proposed in a nonreciprocal Dicke model, with quantum-light-mediated nonreciprocal interactions between two spin ensembles Chiacchio23 . Separately, two groups have independently proposed realizing a nonreciprocal SQPT by exploiting the Sagnac-Fizeau shift in a spinning cavity Zhu24 ; Xu24 . However, the key elements of these proposals-specifically, the required nonreciprocal interactions between two spin ensembles Chiacchio23 and the Sagnac-Fizeau shift in a spinning resonator Zhu24 ; Xu24 -present major challenges for existing experimental quantum platforms that have demonstrated the SQPT Baumann10 ; Baumann11 ; Brennecke13 ; Baden14 ; Zhang21Chen ; Wu23 ; Feng15 ; Zheng23 ; Ning24 ; Cai21 ; Zhao25 ; Chen21 ; Wu24 . It is therefore imperative to explore new theoretical mechanisms for realizing a nonreciprocal SQPT.
Owing to their excellent tunability and design flexibility, cavity magnonic systems have emerged as well-suited platforms for exploring novel physical phenomena Soykal10 ; Huebl13 ; Tabuchi14 ; Zhang14 ; Goryachev14 ; Zhang15 ; Zhang15Zou ; Bai15 ; Lachance-Quirion19 ; Rameshti22 ; Yuan22 . In Ref. Wang16 , frequency shifts of both the magnon and cavity modes induced by the magnon Kerr effect (MKE) were experimentally observed by directly driving a yttrium iron garnet (YIG) sphere within a microwave cavity. The MKE, which originates from magnetocrystalline anisotropy in YIG, describes magnon-magnon interactions. It can lead to multistability of cavity magnon polaritons Wang18 ; Zhang19 ; Nair20 ; Bi21 ; Shen21 ; Shen22 , magnon-photon entanglement Zhang19Scully ; Yang21Jin , long-distance spin-spin coupling Xiong22Tian , and unidirectional microwave transmission Kong19 ; Ullah24 ; Miao24 , among other effects. A notable feature of the MKE is that the Kerr coefficient depends on the direction of the bias magnetic field applied to the YIG sphere: the Kerr coefficient is positive when the field is aligned with the crystallographic axis [100], and negative when aligned with the axis [110] Wang18 ; Zhang19 . It is this directional dependence that makes cavity magnonic systems a promising platform for studying nonreciprocal quantum effects, such as nonreciprocal entanglement Chen23 ; Chen24 ; Ahmed24 ; Kong24 , nonreciprocal steering Kong25 , nonreciprocal magnon blockade Zhang24 , and nonreciprocal mechanical squeezing Tong25 .
In this work, we propose a scheme for realizing a nonreciprocal SQPT induced by MKE in a cavity magnonic system. The system consists of a magnon mode in a YIG sphere coupled to a cavity mode via the magnetic-dipole interaction, where the cavity is subject to a parametric drive, and the magnon mode exhibits the MKE. By analyzing the stability of the solutions for the magnon occupation, we obtain the steady-state phase diagram, which comprises three distinct phases: normal phase, superradiant phase and bistable phase. The phase diagram differs significantly when the bias magnetic field is aligned along the [100] or [110] crystallographic axis of the YIG sphere, corresponding to a positive ($K>0$) or negative ($K<0$) Kerr coefficient, respectively Wang18 ; Zhang19 . We further examine the steady-state magnon number and its fluctuations as functions of the parametric drive strength, revealing that the SQPT becomes nonreciprocal, with different critical thresholds for $K>0$ and $K<0$. To quantify this behavior, we introduce a bidirectional contrast ratio that characterizes the degree of nonreciprocity in the SQPT. Our work provides an alternative approach for realizing a nonreciprocal SQPT Chiacchio23 ; Zhu24 ; Xu24 and offers useful guidance for the experimental realization of this phenomenon in the future.
The remainder of this paper is structured as follows. In Sec. II, we introduce the proposed cavity magnonic system and derive the steady-state solutions for the magnon occupation. Section III presents the steady-state phase diagram based on a stability analysis of these solutions, followed by a detailed investigation of the MKE-induced nonreciprocal SQPT. Finally, a brief summary is given in Sec. IV.
Figure 1: Schematic diagram of the proposed cavity magnonic system. The system comprises a parametrically driven microwave cavity and a YIG sphere magnetized to saturation by a bias magnetic field $B_{0}$. The magnetic field is aligned along either the crystallographic axis [100] or [110] (see the red-arrowed lines) of the YIG sphere.
II Model
As depicted in Fig. 1, the considered cavity magnonic system consists of a microwave cavity and a millimeter-sized YIG sphere, where the cavity is pumped by a parametric drive, and the YIG sphere is magnetized to saturation by a bias static magnetic field. The YIG sphere supports a Kittle mode (with frequency $\omega_{m}$), which is coupled to the cavity mode (with frequency $\omega_{a}$) via the collective magnetic-dipole interaction. Taking into account both the system’s dissipation and the interaction among magnons (i.e., MKE), the parametrically driven cavity magnonic system can be described by the following effective non-Hermitian Hamiltonian (see Appendix A):
$H_{\rm eff}$ $=$ $(\Delta_{a}-i\kappa_{a})a^{{\dagger}}a+(\Delta_{m}-i\gamma_{m})m^{{\dagger}}m+\frac{K}{2}m^{{\dagger}}m^{{\dagger}}mm$ (1)
$+g_{m}(a^{{\dagger}}m+am^{{\dagger}})+\frac{\Omega}{2}(a^{{\dagger}}a^{{\dagger}}+aa),$
where $a^{{\dagger}}$ and $a$ ( $m^{{\dagger}}$ and $m$) are the creation and annihilation operators of the cavity mode (magnon mode), $\Delta_{a(m)}=\omega_{a(m)}-\omega_{d}/2$ $(>0)$ is the frequency detuning of the cavity mode (magnon mode) from the drive field, $\kappa_{a}$ ($\gamma_{m}$) is the decay rate of the cavity mode (magnon mode), $K$ is the Kerr coefficient, $g_{m}$ is the coupling strength between cavity mode and magnon mode, and $\omega_{d}$ ($\Omega$) is the frequency (strength) of parametric drive. Here the MKE results from the magnetocrystalline anisotropy of the YIG sphere, and the sign of Kerr coefficient $K$ can be controlled via adjusting the direction of the bias magnetic field Wang18 ; Zhang19 . When the bias magnetic field is aligned along the crystallographic axis [100], the Kerr coefficient $K$ is positive ($K>0$). In contrast, $K$ becomes negative ($K<0$) for aligning the bias magnetic field along the crystallographic axis [110]. The dependence of $K$ on the direction of the bias magnetic field is the origin of the nonreciprocal SQPT (cf. Sec. III). The effective Hamiltonian $H_{\rm eff}$ maintains invariance under the transformation $a\rightarrow-a$ and $b\rightarrow-b$, demonstrating parity symmetry Emary03PRL ; Emary03PRE . Defining the parity operator $\prod=\exp[i\pi(a^{\dagger}a+b^{\dagger}b)]$, this symmetry can be expressed as the commutation relation $H_{\rm eff},\prod]=0$. In our system, the spontaneous breaking of parity symmetry occurs when the cavity magnonic system undergoes a phase transition from the normal phase to superradiant phase.
Table 1: The parameter conditions (PCs) for $|M|^{2}_{\pm}>0$ in both cases of $K>0$ and $K<0$, where the solution $|M|^{2}=|M|^{2}_{-}$ ($|M|^{2}=|M|^{2}_{+}$) is unstable for all parameter values in the case of $K>0$ ($K<0$) and not considered here.
nonnegativity constraint PCs for $K>0$ PCs for $K<0$
$|M|^{2}_{+}>0$ $\Omega>\Omega_{1}$ for $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}>\xi$ trivial solution
$|M|^{2}_{-}>0$ trivial solution $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{1}$ for $\Delta_{m}/\Delta_{a}>\xi$
The dynamics of the cavity magnonic system can be described through the following quantum Langevin equations Walls94 :
$\dot{a}$ $=$ $-i(\Delta_{a}-i\kappa_{a})a-ig_{m}m-i\Omega a^{{\dagger}}+\sqrt{2\kappa_{a}}\,a_{\rm in},$
$\dot{m}$ $=$ $-i(\Delta_{m}-i\gamma_{m})m-iKm^{{\dagger}}mm-ig_{m}a+\sqrt{2\gamma_{m}}\,m_{\rm in},$ (2)
where $o_{\rm in}$ is the input noise operator for the mode $o$ (o=a,mo=a,\,m), satisfying $\langle o_{\rm in}\rangle=0$. Under the Markovian approximation, the noise operator $o_{\rm in}$ is characterized by the following two-time correlation functions: $\langle o_{\rm in}(t)o_{\rm in}^{\dagger}(t^{\prime})\rangle=[\overline{n}_{o}(\omega_{o})+1]\delta(t-t^{\prime})$ and $\langle o_{\rm in}^{\dagger}(t)o_{\rm in}(t^{\prime})\rangle=\overline{n}_{o}(\omega_{o})\delta(t-t^{\prime})$ with $\overline{n}_{o}(\omega_{o})=[\exp(\hbar\omega_{o}/{k_{B}T})-1]^{-1}$ denoting the equilibrium mean thermal excitation number of the mode $o$, where $k_{B}$ is the Boltzmann constant and $T$ is the bath temperature. To analyze the steady-state behavior of the cavity magnonic system, we perform a linearization of Eq. (II) by decomposing the operators as $a=A+\delta a$ and $m=M+\delta m$, where AA ($M$) represents the expectation value of operator $a$ ($m$), and $\delta a$ ($\delta m$) denotes the quantum fluctuation with $\langle\delta a\rangle=\langle\delta m\rangle=0$. Following from Eq. (II), the expectation values $A$ and $M$ satisfy
$\dot{A}$ $=$ $-i(\Delta_{a}-i\kappa_{a})A-ig_{m}M-i\Omega A^{*},$
$\dot{M}$ $=$ $-i(\Delta_{m}+K|M|^{2}-i\gamma_{m})M-ig_{m}A,$ (3)
and the fluctuation operators $\delta a$ and $\delta m$ obey
$\delta\dot{a}$ $=$ $-i(\Delta_{a}-i\kappa_{a})\delta a-ig_{m}\delta m-i\Omega\delta a^{{\dagger}}+\sqrt{2\kappa_{a}}a_{\rm{in}},$
$\delta\dot{m}$ $=$ $-i(\widetilde{\Delta}_{m}-i\gamma_{m})\delta m-ig_{m}\delta a-iF\delta m^{{\dagger}}+\sqrt{2\gamma_{m}}m_{\rm{in}},$ (4)
where $\widetilde{\Delta}_{m}=\Delta_{m}+2K|M|^{2}$ and $F=KM^{2}$. For obtaining the linear fluctuation equation (II), all small high-order fluctuation terms have been neglected. The steady-state solutions for $|M|^{2}$ and $|A|^{2}$, obtained from Eq. (II) with $\dot{A}=\dot{M}=0$, exhibit three branches due to the MKE:
$|M|^{2}_{0}=0,~~~|M|^{2}_{\pm}=\Big(-\Delta^{\prime}_{m}\pm\sqrt{\eta^{2}G^{2}-{\gamma^{\prime}_{m}}^{2}}\Big)/K,$ (5)
and
$|A|^{2}_{0}=0,~~~|A|^{2}_{\pm}=\left[\big(\Delta_{m}+K|M|_{\pm}^{2}\big)^{2}+\gamma_{m}^{2}\right]|M|_{\pm}^{2}/g_{m}^{2},$ (6)
where $\Delta^{\prime}_{m}=\Delta_{m}-\eta\Delta_{a}$, $\gamma^{\prime}_{m}=\gamma_{m}+\eta\kappa_{a}$ and $\eta=g_{m}^{2}/(\Delta_{a}^{2}+\kappa_{a}^{2}-\Omega^{2})$. Obviously, the photon number $|A|^{2}$ is dependent on the magnon number $|M|^{2}$.
In this study, we choose the magnon number $|M|^{2}$ as the order parameter for the SQPT. The nonnegativity constraint on the magnon number $|M|^{2}$ restricts the valid solutions to $|M|^{2}_{\pm}\geq 0$. Table 1 summarizes the conditions for $|M|^{2}_{+}>0$ when $K>0$ and $|M|^{2}_{-}>0$ when $K<0$, where we exclude the trivial solutions $|M|^{2}_{-}$ when $K>0$ and $|M|^{2}_{+}$ when $K<0$ due to their instability across all parameters (see Fig. 2 and related discussions). For $K>0$, the constraint $|M|^{2}_{+}>0$ requires $\Omega>\Omega_{1}$ when $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}>\xi$, with the critical ratio $\xi$ defined as
$\xi=\frac{2\gamma_{m}^{2}}{\sqrt{4(\Delta_{a}^{2}+\kappa_{a}^{2})\gamma_{m}^{2}+(4\kappa_{a}\gamma_{m}+g_{m}^{2})g_{m}^{2}}-(2\kappa_{a}\gamma_{m}+g_{m}^{2})}.$ (7)
The critical drive strengths $\Omega_{1}$ and $\Omega_{2}$ are respectively given by
$\Omega_{1}=\sqrt{\Delta_{a}^{2}+\kappa_{a}^{2}+(4\kappa_{a}\gamma_{m}+g_{m}^{2})g_{m}^{2}/4\gamma_{m}^{2}}-g_{m}^{2}/2\gamma_{m},$ (8)
and
$\Omega_{2}=\sqrt{\Delta_{a}^{2}+\kappa_{a}^{2}-(2\Delta_{a}\Delta_{m}-2\kappa_{a}\gamma_{m}-g_{m}^{2})g_{m}^{2}/(\Delta_{m}^{2}+\gamma_{m}^{2})},$~~~ (9)
which determined the boundaries between different phases in the steady-state phase diagram of the cavity magnonic system (see Fig. 2). In particular, $\Omega_{1}=\Omega_{2}$ at $\Delta_{m}/\Delta_{a}=\xi$. For $K<0$, the constraint $|M|^{2}_{-}>0$ is satisfied when $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{1}$ for $\Delta_{m}/\Delta_{a}>\xi$.
Figure 2: Steady-state phase diagram of the cavity magnonic system versus the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$, where (a) $K>0$ and (b) $K<0$. NP, SP, and BP represent the normal phase, superradiant phase, and bistable phase, respectively. The white background in the upper-right corner of panel (b) indicates the unstable region. The vertical blue solid line marks the critical drive strength $\Omega=\Omega_{1}$, while the green dashed-dotted curve indicates the critical drive strength $\Omega=\Omega_{2}$. Other parameters are chosen to be $\Delta_{a}/\kappa_{a}=3$, $g_{m}/\kappa_{a}=2.4$, and $\gamma_{m}/\kappa_{a}=1$.
III Nonreciprocal superradiant quantum phase transition
For examining the stability of the solutions for $|M|^{2}$ in Eq. (5), we need to investigate the fluctuations of the system. By introducing the quadrature operators $x_{a}=(\delta a^{{\dagger}}+\delta a)/\sqrt{2}$, $y_{a}=i(\delta a^{{\dagger}}-\delta a)/\sqrt{2}$, $x_{m}=(\delta m^{{\dagger}}+\delta m)/\sqrt{2}$, and $y_{m}=i(\delta m^{{\dagger}}-\delta m)/\sqrt{2}$, we recast Eq. (II) and its complex conjugate expression into the compact matrix form:
$\dot{u}(t)=\mathbf{\Lambda}\,u(t)+f_{\rm in}(t),$ (10)
where $u(t)=[x_{a}(t),\,y_{a}(t),\,x_{m}(t),\,y_{m}(t)]^{T}$ represents the vector of quadrature components, $f_{\rm in}(t)=[\sqrt{2\kappa_{a}}\,x^{\rm(in)}_{a},\,\sqrt{2\kappa_{a}}\,y^{\rm(in)}_{a},\,\sqrt{2\gamma_{m}}\,x^{\rm(in)}_{m},\,\sqrt{2\gamma_{m}}\,y^{\rm(in)}_{m}]^{T}$ denotes the vector of noise quadratures, defined as $x^{\rm(in)}_{a}=(a_{\rm in}^{{\dagger}}+a_{\rm in})/\sqrt{2}$, $y^{\rm(in)}_{a}=i(a_{\rm in}^{{\dagger}}-a_{\rm in})/\sqrt{2}$, $x^{\rm(in)}_{m}=(m_{\rm in}^{{\dagger}}+m_{\rm in})/\sqrt{2}$, and $y^{\rm(in)}_{m}=i(m_{\rm in}^{{\dagger}}-m_{\rm in})/\sqrt{2}$, and the drift matrix $\Lambda$ is given by
$\mathbf{\Lambda}=\left(\begin{array}[]{cccc}-\kappa_{a}&\Delta_{a}-\Omega&0&g_{m}\\ -\Delta_{a}-\Omega&-\kappa_{a}&-g_{m}&0\\ 0&g_{m}&-\gamma_{m}+{\rm Im}[F]&\widetilde{\Delta}_{m}-{\rm Re}[F]\\ -g_{m}&0&-\widetilde{\Delta}_{m}-{\rm Re}[F]&-\gamma_{m}-{\rm Im}[F]\\ \end{array}\right).$ (11)
The stability of solutions $|M|_{0}^{2}$ and $|M|_{\pm}^{2}$ is governed by the spectral properties of the corresponding drift matrix $\mathbf{\Lambda}$ Gradshteyn80 . A given solution is stable only if all eigenvalues $\lambda_{k}$ of matrix $\mathbf{\Lambda}$ satisfy ${\rm Re}(\lambda_{k})<0$ for k=1, 2, 3, 4k=1,\,2,\,3,\,4; otherwise, it is unstable.
Table 2: Stable solutions ($S$s) in each phase for both cases of $K>0$ and $K<0$.
phase type SSs for $K>0$ SSs for $K<0$
normal phase $|M|^{2}=0$ $|M|^{2}=0$
superradiant phase $|M|^{2}=|M|^{2}_{+}$ $|M|^{2}=|M|^{2}_{-}$
bistable phase $|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{+}$ $|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{-}$
Based on the stability analysis of the solutions in Eq. (5), we plot the steady-state phase diagram of the system as a function of the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$ in Fig. 2. Note that the white background in the upper-right corner of Fig. 2(b) indicates the unstable regime, where all three solutions for $|M|^{2}$ are unstable. Figures 2(a) and 2(b) correspond to $K>0$ and $K<0$, respectively, which can be realized by aligning the bias magnetic field parallel to the [100] and [110] crystallographic axes of the YIG sphere Wang18 ; Zhang19 . The difference between these two panels demonstrates the nonreciprocity of the SQPT. The phase diagrams exhibits three distinct phases: normal phase (light blue area), superradiant phase (light green area) and bistable phase (light orange area), where the boundaries among them are determined by the vertical blue solid line $\Omega=\Omega_{1}$ given in Eq. (8)] and the blue dash-dotted curve $\Omega=\Omega_{2}$ given in Eq. (9)]. Table 2 summarizes the stable solutions in each phase. In the normal phase, the solution $|M|^{2}=0$ is stable for both $K>0$ and $K<0$, corresponding to the vacuum state. In the superradiant phase, the solution $|M|^{2}=|M|^{2}_{+}$ for $K>0$ ($|M|^{2}=|M|^{2}_{-}$ for $K<0$) becomes stable, and the system exhibits macroscopic magnon excitations. In the bistable phase, both $|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{+}$ for $K>0$ ($|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{-}$ for $K<0$) are stable, and the system’s initial state determines whether macroscopic magnon excitations occur Zhang21-Chen . For the parameters used in Fig. 2, the critical ratio $\xi$ from Eq. (7) is $\xi=0.976$. When $\Delta_{m}/\Delta_{a}<\xi$, the boundary between normal phase and bistable phase (normal phase and superradiant phase) is given by $\Omega=\Omega_{1}$ ($\Omega=\Omega_{2}$) for $K>0$ ($K<0$). Conversely, when $\Delta_{m}/\Delta_{a}>\xi$, the boundary between normal phase and superradiant phase (normal phase and bistable phase) is given by $\Omega=\Omega_{2}$ ($\Omega=\Omega_{1}$) for $K>0$ ($K<0$). Regardless of the value of $\Delta_{m}/\Delta_{a}$ and the sign of $K$, the curve $\Omega=\Omega_{2}$ universally defines the boundary between bistable phase and superradiant phase.
Figure 3: (a),(b) Scaled steady-state magnon numbers $|M|^{2}/|(\gamma_{m}/K)|$ and (c),(d) magnon number fluctuations $\lg(\langle\delta m^{\dagger}\delta m\rangle+1)$ versus the normalized drive strength $\Omega/\kappa_{a}$, where $T=0$. Yellow shaded regions mark the nonreciprocal domains of SQPT, where $|M|^{2}(K>0)\neq|M|^{2}(K<0)$. The red solid curves correspond to $K>0$, and the blue dashed curves correspond to $K<0$. The detuning ratios are $\Delta_{m}/\Delta_{a}=1.3$ and $\Delta_{m}/\Delta_{a}=0.8$ in panels (a),(c) and (b),(d), respectively. Other parameters are the same as those in Fig. 2.
In Figs. 3(a), we displays the scaled steady-state magnon number $|M|^{2}/|(\gamma_{m}/K)|$ as a function of the normalized drive strength $\Omega/\kappa_{a}$ for a fixed detuning ratio $\Delta_{m}/\Delta_{a}=1.3$ ($\xi=0.976$). Yellow shaded regions indicate the nonreciprocal domains of SQPT, where $|M|^{2}(K>0)\neq|M|^{2}(K<0)$. In the case of $K<0$, the magnon number exhibits a discontinuous jump from zero to a finite value at the critical point $\Omega/\kappa_{a}=2.025$ as the normalized drive strength $\Omega/\kappa_{a}$ increases across this threshold (blue dashed curve), signifying a second-order SQPT from the normal phase to the superradiant phase. However, for $K>0$, the transition becomes first-order (i.e., a continuous SQPT), and the critical point shifts to $\Omega/\kappa_{a}=2.108$ (red solid curve). Figure 3(b) shows $|M|^{2}/|(\gamma_{m}/K)|$ versus $\Omega/\kappa_{a}$ by taking $\Delta_{m}/\Delta_{a}=0.8$ ($\xi=0.976$), where nonreciprocal SQPT is also evident. Unlike the case with $\Delta_{m}/\Delta_{a}>\xi$ [cf. Figs. 3(a) and 3(b)], the system now undergoes a second-order SQPT from the normal phase to the superradiant phase at $\Omega/\kappa_{a}=2.084$ for $K<0$ (blue dashed curve), while the transition remains first-order at $\Omega/\kappa_{a}=2.025$ if $K>0$ (red solid curve).
In order to quantitatively characterize the nonreciprocity in the SQPT, we introduce the bidirectional contrast ratio:
$\mathcal{I}=\left\{\begin{aligned} &0,&&|M|^{2}(K>0)=|M|^{2}(K<0),\\ &\left|\frac{|M|^{2}(K>0)-|M|^{2}(K<0)}{|M|^{2}(K>0)+|M|^{2}(K<0)}\right|,&&|M|^{2}(K>0)\neq|M|^{2}(K<0),\end{aligned}\right$. (12)
which satisfies $0\leq\mathcal{I}\leq 1$. This parameter quantifies the degree of nonreciprocity in the SQPT: a nonzero $\mathcal{I}$ signals the emergence of nonreciprocal behavior. The minimum value $\mathcal{I}_{\rm min}=0$ corresponds to a reciprocal SQPT, while the maximum value $\mathcal{I}_{\rm max}=1$ represents an ideal nonreciprocal SQPT. In Fig. 4, we show the bidirectional contrast ratio $\mathcal{I}$ as a function of the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$. The white region in the upper-right corner corresponds to the unstable regime identified in Fig. 2(b). The vertical blue solid line (red dashed-dotted curve) marks the critical drive strength $\Omega=\Omega_{1}=2.025$ ($\Omega=\Omega_{2}$). The bidirectional contrast ratio $\mathcal{I}$ attains its minimum value $\mathcal{I}_{\rm min}=0$ in the black region ($\Omega<\Omega_{1}$), where the system remains in the normal phase for both $K>0$ and $K<0$, indicating reciprocal SQPT. Notably, two red regions appear for $\Omega_{1}<\Omega<\Omega_{2}$, where $\mathcal{I}$ reaches its maximum value $\mathcal{I}_{\rm max}=1$, corresponding to an ideal nonreciprocal SQPT. In the lower (upper) red region, the system is in the superradiant phase (normal phase) for $K>0$ but remains in the normal phase (superradiant phase) for $K<0$. In the remaining region ($\Omega>\Omega_{2}$), the bidirectional contrast ratio lies between 0 and 1 (i.e., $0<\mathcal{I}<1$). Here, the system is in the superradiant phase for both $K>0$ and $K<0$, but the order parameters differ: $|M|^{2}(K>0)=|M|^{2}_{+}\neq|M|^{2}(K<0)=|M|^{2}_{-}$.
Figure 4: Bidirectional contrast ratio $\mathcal{I}$ as a function of the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$. The white background in the upper-right corner corresponds to the unstable region in Fig. 2(b). The vertical blue solid line marks the critical drive $\Omega=\Omega_{1}$, and the red dashed-dotted curve indicates $\Omega=\Omega_{2}$. Other parameters are the same as those in Fig. 2.
To show more characteristics of the nonreciprocal SQPT, we examine the magnon number fluctuations $\langle\delta m^{\dagger}\delta m\rangle$ on top of the stable mean-field solutions for $|M|^{2}$. The quantum fluctuations of the cavity magnonic system are described by a $4\times 4$ covariance matrix $\mathcal{V}(t)$, with elements defined as $\mathcal{V}_{\rm ij}(t)=\langle u_{i}(t)u_{j}(t^{\prime})+u_{j}(t^{\prime})u_{i}(t)\rangle/2$ (i,j=1, 2, 3, 4)(i,\,j=1,\,2,\,3,\,4). In our model, the magnon number fluctuations $\langle\delta m^{\dagger}\delta m\rangle=\left[(V_{33}+V_{44})-1\right]/2$ are related to the diagonal elements of the steady-state covariance matrix $\mathcal{V}(+\infty)$ Zhu20 , which can be straightforwardly obtained by solving the so-called Lyapunov equation
$\Lambda\mathcal{V}+\mathcal{V}\Lambda^{T}=-\mathcal{D},$ (13)
where the diffusion matrix $\mathcal{D}$, determined by $\mathcal{D}_{\rm ij}\delta(t-t^{\prime})=\langle f_{{\rm in},i}(t)f_{{\rm in},j}(t^{\prime})+f_{{\rm in},j}(t^{\prime})f_{{\rm in},i}(t)\rangle/2$, takes the form $\mathcal{D}={\rm diag}\left[(2\bar{n}_{a}+1)\kappa_{a},(2\bar{n}_{a}+1)\kappa_{a},(2\bar{n}_{m}+1)\gamma_{m},(2\bar{n}_{m}+1)\gamma_{m}\right]$. For the nonreciprocal behavior of the SQPT shown in Figs. 3(a) and 3(b), the corresponding magnon number fluctuations are presented in Figs. 3(c) and 3(d), respectively. For clarity, we plot $\lg(\langle\delta m^{\dagger}\delta m\rangle+1)$ as a function of the normalized drive strength $\Omega/\kappa_{a}$. Here, the red solid curves correspond to $K>0$, and the blue dashed curves to $K<0$. The magnon number fluctuations exhibit clear nonreciprocity, consistent with the nonreciprocal behavior of SQPT. In all cases, the fluctuations diverge near the critical point of the SQPT but approach zero far from the critical point [cf. Figs. 3(a) and 3(b)].
IV Conclusion
In summary, we have demonstrated that the MKE in a cavity magnonic system offers a viable pathway to achieve a nonreciprocal SQPT Chiacchio23 ; Zhu24 ; Xu24 . The proposed setup, comprising a YIG sphere coupled to a microwave cavity, exhibits markedly different phase diagrams under positive and negative Kerr coefficients-controlled by orienting the bias magnetic field along different crystallographic axes of the YG sphere Wang18 ; Zhang19 . These distinct phase boundaries directly lead to nonreciprocal critical behavior, as confirmed by analyzing the magnon occupation and its fluctuations around the transition. The introduction of a bidirectional contrast ratio further allows quantitative characterization of the nonreciprocity. Our results establish cavity magnonics as a promising platform for exploring nonreciprocal SQPT and may inform future designs of quantum devices requiring directional control.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grants No. 12205069 and No. 12204139), the National Key Research and Development Program of China (Grant No. 2024YFA1408900), the Natural Science Foundation of Zhejiang Province (Grant No. LY24A040004), Zhejiang Province Key R&D Program of China (Grant No. 2025C01028), and Shenzhen International Quantum Academy (Grant No. SIQA2024KFKT010).
Appendix A Derivation of the effective non-Hermitian Hamiltonian in Eq. (1)
As shown in Fig. 1, the magnon mode in the YIG sphere is coupled to the cavity mode. Including both the parametric drive and the MKE, the total Hamiltonian of the system is given by
$H$ $=$ $\omega_{a}a^{{\dagger}}a+\omega_{m}m^{{\dagger}}m+\frac{K}{2}m^{{\dagger}}m^{{\dagger}}mm+g_{m}(a^{{\dagger}}m+am^{{\dagger}})$ (14)
$+\frac{\Omega}{2}(a^{{\dagger}}a^{{\dagger}}e^{-i\omega_{d}t}+aae^{i\omega_{d}t}),$
where $a^{{\dagger}}$ and $m^{{\dagger}}$ ($a$ and $m$) are the creation (annihilation) operators for the cavity and magnon modes with frequencies $\omega_{a}$ and $\omega_{m}$, respectively, $K$ is the Kerr coefficient, $g_{m}$ is the coupling strength between cavity and magnon modes, and $\Omega$ ($\omega_{d}$) is the strength (frequency) of parametric drive. In the rotating frame at $\omega_{d}/2$, the Hamiltonian in Eq. (14) becomes
$H$ $=$ $\Delta_{a}a^{{\dagger}}a+\Delta_{m}m^{{\dagger}}m+\frac{K}{2}m^{{\dagger}}m^{{\dagger}}mm+g_{m}(a^{{\dagger}}m+am^{{\dagger}})$ (15)
$+\frac{\Omega}{2}(a^{{\dagger}}a^{{\dagger}}+aa),$
where $\Delta_{a}=\omega_{a}-\omega_{d}/2$ and $\Delta_{m}=\omega_{m}-\omega_{d}/2$ represent the detunings of the cavity and magnon modes from the parametric drive, respectively. With the Heisenberg-Langevin approach Walls94 , the quantum dynamics of the system is governed by
$\dot{a}$ $=$ $-i(\Delta_{a}-i\kappa_{a})a-ig_{m}m-i\Omega a^{{\dagger}}+\sqrt{2\kappa_{a}}\,a_{\rm in},$
$\dot{m}$ $=$ $-i(\Delta_{m}-i\gamma_{m})m-iKm^{{\dagger}}mm-ig_{m}a+\sqrt{2\gamma_{m}}\,m_{\rm in},$~~~ (16)
which is Eq. (II) in the main text. Here, $\kappa_{a}$ ($\gamma_{m}$) denotes the decay rate of the cavity mode (magnon mode), and $a_{\rm in}$ ($m_{\rm in}$) is the corresponding input noise operator. Rewriting Eq. (A) in the form
$\dot{a}$ $=$ $-i[a,H_{\rm eff}]+\sqrt{2\kappa_{a}}\,a_{\rm in},$
$\dot{m}$ $=$ $-i[m,H_{\rm eff}]+\sqrt{2\gamma_{m}}\,m_{\rm in},$ (17)
yields the effective non-Hermitian Hamiltonian $H_{\rm eff}$ presented in Eq. (1).